# Disentangling the Complex Multiplexed DIA Spectra in De Novo Peptide Sequencing

### v5

@@ 3 Discussion
This study introduces a robust and highly accurate method for de novo peptide sequencing within the DIA setting, offering substantial improvements over existing approaches. By incorporating dilated convolutional neural networks, FlashTtention-2, RoPE, as well as coelution-aware pretraining, our model is adept at handling the complex, highly-multiplexed and noisy spectral data inherent to DIA experiments. The use of these advanced techniques allows our approach to effectively capture intricate spectral patterns, enabling reliable peptide sequence predictions even in challenging conditions.
@@
Our comprehensive experiments demonstrate the superior performance of the proposed method, particularly when applied to data from the Orbitrap Astral mass spectrometer. The Astral’s enhanced fragment ion coverage and consistent fragmentation patterns provide a solid foundation for our de novo sequencing approach, leading to a marked increase in peptide identification accuracy, even in case of high coelution level. In contrast, older-generation mass spectrometers, such as those used in older DDA and DIA experiments, reveal the limitations of existing de novo sequencing methodologies, especially when larger isolation windows are used. Our DDA vs. DIA comparison study highlights these challenges, showing that while DIA outperforms DDA with narrower isolation windows, it begins to lose this advantage as the window size increases. This issue is particularly evident with older-generation instruments, where DIA performance in both database search and de novo sequencing diminishes at larger window sizes. However, the Orbitrap Astral consistently demonstrates DIA’s superiority over DDA, facilitated by the narrow isolation windows enabled by this instrument, underscoring the pivotal role that modern mass spectrometers play in advancing peptide sequencing capabilities. In addition, our ablation study verifies the validity and neccessity of our design.
@@
During our theoretical analysis of de novo performance, we show that the balance between signal enhancement and noise accumulation is critical. In the case of Astral DIA, the substantial increase in informative signal peaks provides a richer dataset for de novo algorithms to process, improving the accuracy and confidence of peptide sequence reconstruction. This outcome underscores the importance of optimizing acquisition parameters to maximize signal quality while managing noise levels.
@@
Our findings suggest that acquisition methods like Astral DIA, which can substantially increase signal peaks while effectively managing noise, are more conducive to de novo peptide sequencing. This has important implications for the development of mass spectrometry techniques and the selection of acquisition parameters in proteomics research. Conversely, older-generation DIA does not offer the same advantage because the modest gain in signal peaks is insufficient to counterbalance the significant escalation in noise peaks. This results in lower confidence in peptide identifications, as reflected by higher p-values.
In conclusion, our study represents an advancement in de novo peptide sequencing within the DIA framework. By integrating advanced computational techniques to tackle the coeluted nature of DIA spectra, and leveraging the advanced capabilities of next-generation mass spectrometers like the Orbitrap Astral, we offer a robust and accurate solution for analyzing complex proteomic data. The findings of this study not only demonstrate the limitations of older-generation instruments in de novo peptide sequencing but also highlight the potential of DIA to provide more comprehensive and precise peptide identification. This work paves the way for future innovations in de novo sequencing, offering new opportunities for discovering novel peptides and advancing our understanding of proteomic diversity.
4 Methods
Our model architecture is an encoder-decoder Transformer operating on the spectrum graph, including spectrum graph construction, dilated CNN (convolutional neural network) layers to process the chromatograms, RoPE (rotary positinal embedding)[40] integration to encode mass differences in the graph, coelution-aware pretraining to provide better understanding of the multiplexed spectra, and the two-stage decoding, where we predict the optimal path through the spectrum graph in stage 1, and refine the path to generate the final peptide sequence with mass tag filling in stage 2, in order to alleviate the sequence memorization problem commonly associated with de novo sequencing algorithms, where the model relies too much on previous seen peptide sequences during training. The model architecture is shown in Figure 9.
Figure 9: The model structure of Our entire workflow. On the top is the optimal path task, generating a series of node indices, which are transformed into the optimal path. The mass values in the optimal path are then translated to the corresponding amino acids when a single match is found. On the button is the sequence generation task. It takes the generated optimal path as input and outputs the amino acid sequence to replace mass tags. In the figure, FFNGLU refers to Feed-Forward Network with Gated Linear Units, commonly used in Transformers and deep learning architectures to enhance expressiveness and efficiency. CNN refers to Convolutional Neural Networks, and RoPE refers to Rotary Positional Embeddings.
4.1 Feature Extraction
@@
The feature construction for the encoder is designed to capture the chromatogram characteristics in DIA, in order to better distinguish signal peaks from coeluting fragment ions and noise. For each PSM (peptide-spectrum match), we collect five neighboring MS2 spectra along the retention time (RT) dimension, centered around the RT peak. This selection is crucial as it differentiates DIA data from DDA. The neighboring spectra provide a chromatogram for each observed peak in the spectrum, and the relationship among these chromatograms improves the robustness of peptide sequence predictions. [29]
For MS1 spectra, alignment with corresponding MS2 spectra is needed. To achieve this, MS1 intensity values are linearly interpolated based on their RT values to match the MS2 spectra exactly, ensuring that the features derived from MS1 and MS2 spectra are directly comparable.
During preprocessing, each spectrum is binned into fixed-size 0.01 Th intervals, enabling the construction of chromatograms for each m/z value. Once the chromatograms are generated through binning, the spectra are reconstructed as a list of (m/z, chromatogram) pairs for neural network processing.
In addition to the primary spectral data, our methodology involves the calculation of eight supplementary features to enrich the feature set used for encoding, adapted from GraphNovo[13] and Novor[41]. These features are derived from each spectrum within the ±$\pm 50$ Th neighboring window and are designed to capture various statistical and intensity-based characteristics of the spectral peaks.
@@
Finally, the spectrum graph is constructed by transforming every peak in the original spectrum from m/z to N-terminal residual masses. Each peak in the original spectrum is converted into 6 peaks in spectrum graph, taking into account 6 types of ions, including $a^{+}$, $a^{2+}$, $b^{+}$, $b^{2+}$, $y^{+}$, and $y^{2+}$. We operate on the N-terminal residual masses, denoted as the graph node mass, following Equation 1, during which C-terminal ions are converted to their N-terminal residual masses by subtracting their C-terminal residual mass from the precursor mass. Note that although we did not specifically include other common ion types like $b-H_{2}O$ or $y-{NH_{3}}$, their peaks are converted into graph node as well. They do not belong to the target sequence, but their graph node can still contribute to our de novo objective by self attention. We provide an example of spectrum graph in Figure 10.
Figure 10: An example of a spectrum graph is shown, where the bottom value on each edge represents the mass difference between nodes, encoded by RoPE, and the top value indicates the corresponding amino acid sequence. Only a subset of nodes and edges is plotted for clarity, whereas, in a complete spectrum graph, all possible forward connections would be present.
$m_{\text{nterm}}=\begin{cases}\left(m/z-m_{\text{proton}}\right)\times c-\sigma,&\text{if ion}\in\{a,b,c\text{-ions}\}\\ m_{\text{precursor}}-\left(m/z-m_{\text{proton}}\right)\times c+\sigma,&\text{if ion}\in\{x,y,z\text{-ions}\}\end{cases}$ (1)
where offset $\sigma$ depends on the ion type, computed as Supporting Table 1; $c$ is the charge of ion; $m_{proton}$ is the mass of proton and $m_{precursor}$ is the precursor mass [13].
4.2 Time-Series Encoder and Spectrum Encoder
In our approach, we employ a dilated convolutional neural network (CNN) to effectively encode the time-series data derived from the chromatogram and the additional spectral features.
Dilated CNNs [42] are a variant of traditional convolutional networks. The primary distinction of dilated convolutions lies in their ability to expand the receptive field without losing resolution or coverage, achieved by inserting gaps (known as dilation) between each element in the convolution kernel. This allows the network to encode time-dependent information in chromatograms while keeping the number of parameters and memory/computational complexity relatively low.
For the chromatogram and its associated spectral features, dilated CNNs are ideally suited. Our model processes inputs structured as $[5,1]$ or $[5,8]$ tensors, where ‘$5$’ represents the number of neighboring spectra considered, ‘$1$’ denotes the chromatogram, and ‘$8$’ reflects the additional computed features. By applying dilated convolutions, the model can integrate both local and broader contextual information across these inputs. The outputs of the dilated CNN are encoded into tensors of size $[hidden\_size]$, where ‘$hidden\_size$’ indicates the dimension of the feature space. This encoding captures information about the temporal dynamics of a specific m/z value.
@@ 4.2 Time-Series Encoder and Spectrum Encoder
In addition, we adopt the FlashAttention 2 [26] implementation of the attention function, allowing us to process very large DIA spectra, with high computation and memory efficiency.
4.3 RoPE Integration
GraphNovo employs the Graphformer[43] graph neural network to encode the spectrum graph, capturing comprehensive information about the spectrum. However, applying this approach to DIA data is impractical due to the large size of DIA spectra, which results from high levels of coelution. To address this challenge, we replace the memory-intensive node and path encoders used in Graphformer with Rotary Position Embedding (RoPE). This step is also neccessary for the adoption of Flash Attention 2, since it does not allow us to operate directly on the attention matrix, which is required by Graphformer.
RoPE is a position encoding technique that introduces rotational invariance by representing positional information in a circular format[40]. In our approach, RoPE encodes the mass differences between graph nodes, allowing the model to learn meaningful mass relationships automatically. This effectively transforms the spectrum graph into a fully-connected graph where edges represent mass differences, enabling the model to identify the most relevant mass differences for the task without the computational overhead of traditional node and path encoders. The adoption of RoPE is advantageous, as it captures the critical information conveyed by the mass difference between pairs of graph nodes, effectively representing the cumulative mass of the amino acids that lie between them.
4.4 Two Stage Decoder
We adopt a two-stage decoding process, similar to GraphNovo[13] to alleviate the sequence memorization issue during training, where the decoder simply remembers seen training sequences, and fail to generalize on unseen sequences. This approach includes:
@@ 4.4 Two Stage Decoder
Readers can refer to Figure 1a of Mao et al. (2023)[13] for a detailed visualization of the two-stage decoding process.
4.5 Coelution-aware Pretraining
@@
In a normal de novo sequencing algorithm, we typically only consider the fragment ions of the de novo target peptide. For instance, in the optimal path task of our program, we predict only the graph nodes (n-terminal masses of peaks) belonging to the de novo target peptide, and the next step we replace these graph nodes with amino acids, resulting in predicted peptide sequence.
However, if we incorporate the information about the fragment ions of other coeluting peptides, we might be able to provide the model broader information about the spectrum, improving de novo performance. More specifically, in traditional de novo algorithms, the fragment ions of other coeluting peptides are treated as noises, but we can give them labels for the model to learn. With such information, the model can make less mistakes distinguishing noise peaks from signal peaks (of the target de novo peptide), since the model knows some noise peak is probably a fragment ion of other coeluting peptide, thus less likely predict it as a signal peak.
To achieve this, we introduce coelution-aware pretraining to our algorithm. We adopt the same Transformer encoder in Section 4.1, without the spectrum graph conversion, i.e. we keep all the original m/z values as the input to the Transformer. After layers of self-attention, we obtain the embedding for each m/z value in the spectrum. These embeddings are trained under the ion type loss, a cross entropy loss, representing the type of fragment ions. 12 types of ions are considered (see Supporting Table 2), plus noise. These labels corresponds to the fragment ion types of all coeluting peptides, not only the target peptide.
The workflow for the pretraining model is illustrated in Figure 11. After pretraining, the trained embeddings are fed to downstream optimal path and sequence generation models as features for the Transformer encoder.
Figure 11: Pretrain model architecture.
4.6 Theoretical Analysis
We implemented simulations to generate synthetic mass spectra reflecting the characteristic signal and noise levels associated with DDA, older-generation DIA, and Astral DIA methods.
@@ 4.6 Theoretical Analysis
    Computation of p-values: We employed the dynamic programming approach described by Noble, et al. [34] to compute the p-values associated with the XCorr scores. This method estimates the distribution of XCorr scores under the null hypothesis and determines the statistical significance of the observed XCorr scores.
@@
    Theoretically, the p-value is inversely related to peptide recall. This inverse relationship arises because, when computing the Sidak-corrected p-value [44], we must adjust for a significantly larger number of peptide hypotheses in de novo sequencing than in database searches. In database searches, only the peptides existing in the database are considered, limiting the number of comparisons. In contrast, de novo sequencing must account for all possible peptide sequences, dramatically increasing the number of hypotheses and necessitating a stricter correction for multiple testing. Consequently, the p-value threshold becomes more stringent in de novo sequencing, potentially reducing peptide recall. The relationship between peptide recall and p-value is given by the following equation:
    $Peptide\ Recall=\mathbb{E}{\left[\frac{\#Denovo\ Success}{\#Database\ Success}\right]}\approx\frac{(1-p)^{\#\ De\ Novo\ Peptides}}{(1-p)^{\#\ Database\ Peptides}}$ (2)
    where #$\#\ De\ Novo\ Success$ and #$\#\ Database\ Success$ are the number of successfully identified peptides by de novo mode/ database search mode, $p$ is the p value, and #$\#\ De\ Novo\ Peptides$ and #$\#\ Database\ Peptides$ refer to the number of peptides with the same precursor mass to compare for de novo or database search, with #$\#\ De\ Novo\ Peptides\gg\#\ Database\ Peptides$
5 Conclusion
In this paper, we present a novel and highly effective method for de novo peptide sequencing in the DIA setting, offering significant improvements over traditional approaches. A key focus of our study is the comparison between DDA and DIA, where we demonstrate that while DIA outperforms DDA with narrow isolation windows, its advantage diminishes with wider windows in older-generation instruments. However, with the Orbitrap Astral, DIA consistently surpasses DDA, highlighting the importance of next-generation mass spectrometers in improving peptide identification. To explain this phenomenon, we propose a simulation based theory.
Our method, specifically designed to tackle the highly mutliplexed DIA data, along with leveraging advanced instrumentation, provides robust performance in challenging proteomic environments. These findings emphasize the value of using DIA with modern technologies for comprehensive peptide detection, paving the way for more reliable and effective de novo sequencing methods in proteomics.
Appendix A Additional Features for Spectrum

### v6

@@ Discussion
This study introduces a robust and highly accurate method for de novo peptide sequencing within the DIA setting, offering substantial improvements over existing approaches. With DIA-oriented design, our model is adept at handling the complex, highly-multiplexed and noisy spectral data inherent to DIA experiments.
@@
Our comprehensive experiments demonstrate the superior performance of the proposed method, particularly when applied to data from the Orbitrap Astral mass spectrometer. The Astral’s enhanced fragment ion coverage and consistent fragmentation patterns provide a solid foundation for our de novo sequencing approach, leading to a marked increase in peptide identification accuracy, even in case of high coelution level. In contrast, older-generation mass spectrometers, such as those used in older DDA and DIA experiments, reveal the limitations of existing de novo sequencing methodologies, especially when larger isolation windows are used. Our DDA vs. DIA comparison study highlights these challenges, showing that while DIA outperforms DDA with narrower isolation windows, it begins to lose this advantage as the window size increases. This issue is particularly evident with older-generation instruments, where DIA performance in both database search and de novo sequencing diminishes at larger window sizes. However, the Orbitrap Astral consistently demonstrates DIA’s superiority over DDA, facilitated by the narrow isolation windows enabled by this instrument, underscoring the pivotal role that modern mass spectrometers play in advancing peptide sequencing capabilities.
Theoretical Analysis
@@
During our theoretical analysis of de novo performance, we show that the balance between signal enhancement and noise accumulation is critical. This outcome underscores the importance of optimizing acquisition parameters to maximize signal quality while managing noise levels. We need to note that de novo peptide sequencing is affected more by the deminishing match quality compared to database search, explaining the wider gap between de novo and database search in wide window settings.
@@
Our findings suggest that acquisition methods like Astral DIA, which can substantially increase signal peaks, despite its higher noise and lower signal-to-noise ratio, are more conducive to de novo peptide sequencing. This has important implications for the development of mass spectrometry techniques and the selection of acquisition parameters in proteomics research. Conversely, older-generation DIA does not offer the same advantage because the modest gain in signal peaks is insufficient to counterbalance the significant escalation in noise peaks, despite their lower signal-to-noise ratio. This results in lower confidence in peptide identifications, as reflected by higher p-values. We can also conclude that identification performance depends on more than just the signal-to-noise ratio, with the absolute strength of signal peaks playing a more decisive role.
Potential Applications of Signal- and Noise-Peak–Based Performance Prediction
We can estimate per-spectrum signal and noise peak numbers without identification by combining an MS1-based purity cue with complementary b/y-ion evidence. Briefly, we measure how much of the local MS1 signal in a DIA frame comes from the dominant precursor (its purity) and use a small, once-calibrated model (driven by precursor intensity, charge, collision energy, resolution, and injection time/AGC) to predict how many fragment peaks that precursor would produce if isolated; scaling this “fragment budget” by purity yields a physics-based signal estimate. In parallel, we scan the MS2 spectrum for complementary b/y pairs whose masses sum to the precursor’s neutral mass (within tolerance) and that co-elute with the precursor in retention time; the union of peaks in validated pairs gives a second, sequence-agnostic signal estimate. Fusing the two gives the signal-peak count; subtracting from the total peaks yields noise.
Compared to performing a small-scale search, this is real-time and engine-independent, provides quality metrics for all spectra (including those that remain unidentified), avoids circularity with scoring/FDR, and scales to rapid “what-if” evaluations of acquisition settings.
This approach leads to many potential applications. For instance, we can perform real-time/adaptive acquisition: sub-second decisions to extend/shorten injection time, adjust collision energy, skip frames, or schedule narrow-window revisits based on the prediction of spectrum quality we receive. Pre-search compute triage could also be achieved by binning spectra by predicted difficulty and route expensive open-mod/de novo only to spectra likely to succeed, improving IDs per CPU/GPU hour.
[…]
DIANovo De Novo Sequencing Model
Our model architecture is based on an encoder–decoder Transformer, augmented with a BERT-style pretraining module. We operate on a spectrum graph representation, in which each spectrum is converted into an aligned format by grouping peaks across adjacent spectra with closely matching $m/z$ values. This representation naturally encodes chromatographic information along the retention time (RT) dimension, which we find to be highly beneficial for improving de novo sequencing performance. During pretraining, the model takes the original spectrum as input and predicts the ion types for all coeluting peptides, rather than focusing solely on the target peptide, as illustrated in Figure 1(b). This strategy could enable the model to better capture the multiplexed nature of DIA spectra. For de novo sequencing, we employ a two-stage decoding framework (Figure 1(a)). In the first stage, the model predicts the optimal path through the spectrum graph, and in the second stage, it refines this path by filling in mass tags to generate the final peptide sequence. This approach mitigates the sequence memorization problem often observed in de novo sequencing algorithms, where models overly rely on peptides frequently seen during training. Additionally, we incorporate rotary positional embeddings (RoPE) to encode the mass differences between spectrum graph nodes, which we find provides superior performance compared to the use of absolute positional embeddings alone. Further details on the model design are provided in Supporting Section A.
Theoretical Analysis
To better understand the statistical behavior of de novo peptide sequencing, we performed controlled simulations to generate synthetic spectra that mimic the signal and noise characteristics of different acquisition methods and test datasets. Signal peaks were sampled from theoretical spectra to represent peptide evidence, while noise peaks were added at levels representative of typical coelution and background noise for each acquisition method. For each simulated spectrum, we computed XCorr scores by cross-correlating the simulated spectra with the corresponding theoretical spectra, followed by significance estimation using a dynamic programming approach for p-value calculation (32). Our analysis further examined the impact of multiple hypothesis testing in de novo sequencing, where the dramatically larger search space compared to database searches necessitates stricter statistical corrections, such as the Šidák adjustment (38). This adjustment increases the stringency of significance thresholds, which in turn reduces peptide recall in de novo sequencing relative to database-driven methods. This theoretical derivation further illustrates that database searches are less impacted by the elevated p-values arising from the degraded signal-to-noise profiles observed when transitioning from DDA to older-generation DIA. This resilience helps database search methods maintain reasonable performance under noisier conditions, whereas de novo sequencing, with its vastly larger hypothesis space, experiences a more pronounced drop in performance.
A full description of the simulation setup, parameter choices, and theoretical derivations supporting this analysis are provided in the Supporting Materials Section B.
Code and Data Availability
[…]
The authors received support from Canadian Network for Research and Innovation in Machining Technology, Natural Sciences and Engineering Research Council of Canada (NSERC Canadian Network for Research and Innovation in Machining Technology), grant number OGP0046506 [Li].
[…]
[34]
Appendix A The DIANovo Model
Feature Extraction
@@ Appendix A The DIANovo Model
The feature construction for the encoder is designed to capture the chromatogram characteristics in DIA, in order to better distinguish signal peaks from coeluting fragment ions and noise. For each PSM (peptide-spectrum match), we collect five neighboring MS2 spectra along the retention time (RT) dimension, centered around the RT peak. The neighboring spectra provide a chromatogram for each observed peak in the spectrum, and the relationship among these chromatograms improves the robustness of peptide sequence predictions. (27). During preprocessing, peaks across neighboring spectra with similar $m/z$ values are grouped together, enabling the construction of chromatograms for each m/z value. This is achieved by binning each spectrum into fixed-size 0.01 Th intervals, and once the chromatograms are generated through binning, the spectra are reconstructed as a list of (m/z, chromatogram) pairs for neural network processing. The binning approach offers greater computational efficiency than methods such as KD-trees, making it particularly well-suited for the characteristics of our DIA data.
For MS1 spectra, alignment with corresponding MS2 spectra is achieved by interpolating MS1 intensity values based on their RT values to match the MS2 spectra, ensuring that the features derived from MS1 and MS2 spectra are directly comparable.
In addition to the primary spectral data, we also included eight supplementary features to enrich the feature set used for encoding, adapted from GraphNovo(14) and Novor(39). These features are derived from each spectrum within the $\pm 50$ Th neighboring window and are designed to capture various statistical and intensity-based characteristics of the spectral peaks.
@@
Finally, the spectrum graph is constructed by transforming every peak in the original spectrum from m/z to N-terminal residual masses. Each peak in the original spectrum is converted into 6 peaks in spectrum graph, taking into account 6 types of ions, including $a^{+}$, $a^{2+}$, $b^{+}$, $b^{2+}$, $y^{+}$, and $y^{2+}$. We operate on the N-terminal residual masses, denoted as the graph node mass, following Equation 1, during which C-terminal ions are converted to their N-terminal residual masses by subtracting their C-terminal residual mass from the precursor mass. Note that although we did not specifically include other common ion types like $b-H_{2}O$ or $y-{NH_{3}}$, their peaks are converted into graph node as well. They do not belong to the target sequence, but their graph node can still contribute to our de novo objective by self attention. We provide an example of spectrum graph in Figure 1(c).
$m_{\text{nterm}}=\begin{cases}\left(m/z-m_{\text{proton}}\right)\times c-\sigma,\\ \hfill\text{if ion}\in\{a,b,c\text{-ions}\}\\ m_{\text{precursor}}-\left(m/z-m_{\text{proton}}\right)\times c+\sigma,\\ \hfill\text{if ion}\in\{x,y,z\text{-ions}\}\end{cases}$ (1)
where offset $\sigma$ depends on the ion type, computed as Supporting Table 1; $c$ is the charge of ion; $m_{proton}$ is the mass of proton and $m_{precursor}$ is the precursor mass (14).
Time-Series Encoder and Spectrum Encoder
In our approach, we employ a dilated convolutional neural network (CNN) to encode the time-series data derived from the chromatogram and the additional spectral features, encoding time-dependent information in chromatograms while keeping the number of parameters and memory/computational complexity relatively low. Our model processes inputs structured as $[5,1]$ or $[5,8]$ tensors, where ‘$5$’ represents the number of neighboring spectra considered, ‘$1$’ denotes the chromatogram, and ‘$8$’ reflects the additional computed features. The outputs of the dilated CNN are encoded into tensors of size [hidden_size][hidden\_size], where ‘hidden_sizehidden\_size’ indicates the dimension of the feature space. This encoding captures information about the temporal dynamics of a specific m/z value.
@@
In addition, we adopt the FlashAttention 2 (40) implementation of the attention function, allowing us to process very large DIA spectra, with high computation and memory efficiency.
RoPE Integration
GraphNovo employs the Graphformer(41) graph neural network to encode the spectrum graph, capturing comprehensive information about the spectrum. However, applying this approach to DIA data is impractical due to the large size of DIA spectra, which results from high levels of coelution. To address this challenge, we replace the memory-intensive node and path encoders used in Graphformer with Rotary Position Embedding (RoPE). This step is also neccessary for the adoption of Flash Attention 2, since it does not allow us to operate directly on the attention matrix, which is required by Graphformer.
RoPE is a position encoding technique that introduces rotational invariance by representing positional information in a circular format(42). In our approach, RoPE encodes the mass differences between graph nodes, allowing the model to learn meaningful mass relationships automatically. This effectively transforms the spectrum graph into a fully-connected graph where edges represent mass differences, enabling the model to identify the most relevant mass differences for the task without the computational overhead of traditional node and path encoders. The adoption of RoPE is advantageous, as it captures the critical information conveyed by the mass difference between pairs of graph nodes, effectively representing the cumulative mass of the amino acids that lie between them.
Two Stage Decoder
We adopt a two-stage decoding process, similar to GraphNovo(14) to alleviate the sequence memorization issue during training, where the decoder simply remembers seen training sequences, and fail to generalize on unseen sequences. This approach includes:
@@
Coelution-aware Pretraining
@@
In a normal de novo sequencing algorithm, we typically only consider the fragment ions of the de novo target peptide. For instance, in the optimal path task of our program, we predict only the graph nodes (n-terminal masses of peaks) belonging to the de novo target peptide, and the next step we replace these graph nodes with amino acids, resulting in predicted peptide sequence. However, if we incorporate the information about the fragment ions of other coeluting peptides, we might be able to provide the model broader information about the spectrum, improving de novo performance. More specifically, in traditional de novo algorithms, the fragment ions of other coeluting peptides are treated as noises, but we can give them labels for the model to learn. With such information, the model can make less mistakes distinguishing noise peaks from signal peaks (of the target de novo peptide), since the model knows some noise peak is probably a fragment ion of other coeluting peptide, thus less likely predict it as a signal peak.
To achieve this, we introduce coelution-aware pretraining to our algorithm. We adopt the same Transformer encoder in Section A, without the spectrum graph conversion, i.e. we keep all the original m/z values as the input to the Transformer. After layers of self-attention, we obtain the embedding for each m/z value in the spectrum. These embeddings are trained under the ion type loss, a cross entropy loss, representing the type of fragment ions. 12 types of ions are considered (see Supporting Table 2), plus noise. These labels corresponds to the fragment ion types of all coeluting peptides, not only the target peptide. After pretraining, the trained embeddings are fed to downstream optimal path and sequence generation models as features for the Transformer encoder.
Appendix B Theoretical Analysis
We implemented simulations to generate synthetic mass spectra reflecting the characteristic signal and noise levels associated with different acquisition methods and datasets.
@@ Appendix B Theoretical Analysis
Computation of p-values: We employed the dynamic programming approach described by Noble, et al. (32) to compute the p-values associated with the XCorr scores. This method estimates the distribution of XCorr scores under the null hypothesis and determines the statistical significance of the observed XCorr scores.
@@
Theoretically, the p-value is inversely related to peptide recall. This inverse relationship arises because, when computing the Sidak-corrected p-value (38), we must adjust for a significantly larger number of peptide hypotheses in de novo sequencing than in database searches. In database searches, only the peptides existing in the database are considered, limiting the number of comparisons. In contrast, de novo sequencing must account for all possible peptide sequences, dramatically increasing the number of hypotheses and necessitating a stricter correction for multiple testing. Consequently, the p-value threshold becomes more stringent in de novo sequencing, potentially reducing peptide recall. The relationship between peptide recall and p-value is given by the following equation:
$Peptide\ Recall$ $=\mathbb{E}{\left[\frac{\#Denovo\ Success}{\#Database\ Success}\right]}$ (2)
$\approx\frac{(1-p)^{\#\ De\ Novo\ Peptides}}{(1-p)^{\#\ Database\ Peptides}}$
where #DeNovoSuccess\#\ De\ Novo\ Success and #DatabaseSuccess\#\ Database\ Success are the number of successfully identified peptides by de novo mode/ database search mode, $p$ is the p value, and #DeNovoPeptides\#\ De\ Novo\ Peptides and #DatabasePeptides\#\ Database\ Peptides refer to the number of peptides with the same precursor mass to compare for de novo or database search, with $De\ Novo\ Peptides\gg\#\ Database\ Peptides$
Additionally, in peptide database search, the true positive is also inversely related to the p-value, as evidenced by the probability of successful database identification computed as
$DB\ Identification\ Prob$ $=\mathbb{E}{\left[\frac{\#Database\ Success}{\#Total}\right]}$ (3)
$\approx(1-p)^{\#\ Database\ Peptides}$
Although both de novo peptide recall and database-based identification probability exhibit an inverse relationship with the corresponding p-values, the impact of increasing p-value is more pronounced in the de novo approach. This trend suggests that de novo identification methods are more sensitive to statistical confidence levels compared to database search strategies, due to their reliance on sequence inference without prior knowledge.
Appendix C Additional Features for Spectrum